# Group-robust Machine Unlearning

### v1

@@ 4 Experiments
This section describes the experimental protocol (Sec. 3.1) and compares MIU with multiple methods in group robust unlearning (Sec. 4.2). We then compare existing approaches and ours when varying the unlearning ratio of the forget set dominant group (Sec. 4.3) and when sampling the forget set from multiple groups (Sec. 4.4). Finally, Sec. 4.5 shows a complete ablation study of MIU’s components.
@@ 4.1 Experimental protocol
Datasets. We follow established works in group-robust optimization [42, 21, 30, 38, 37] and conduct experiments on CelebA [31] and Waterbirds [42] datasets, setting attractive and male as the target and protected attributes for CelebA [31], while setting waterbirds and land as the target and sensitive attributes for Waterbirds [42]. We additionally experiment with FairFace [24], given its numerous annotations, using age as the downstream task and randomly choosing the class 20-29 and the ethnicity afro-american as target and protected attributes, respectively. Unless stated otherwise, forget-set images are sampled from groups described by the above target-sensitive attribute pairs. Moreover, unless stated otherwise, the forget set simulates the worst-case scenario where a single group is responsible for the unlearning request, leading to a high forget distribution imbalance. After unlearning the model must have unlearned the forget data and maintained its original robustness.
@@
Baselines. We compare MIU against three state-of-the-art machine unlearning approaches. L1-sparse [23] forgets sensitive information by fine-tuning the original model on the retain set with a sparsity regularization term. SalUn [13] proposes a saliency-based unlearning that forgets data via random labeling. SCRUB [28] minimizes the divergence between the unlearned and original model on the retain set while maximizing it on the unlearning data. Following previous works [28, 23], we report Pretrain, and Retrain, which are computed by fine-tuning an ImageNet [41] pre-trained ResNet-18 [17]. We also report Retrain + group-DRO [42] to validate the proposed Retrain + reweight. Further details are in the Appendix.
@@
Metrics. To evaluate unlearning and group-robustness, we rely on six different metrics. The first three are, retain (RA), forget (UA), and test (TA) accuracy. We also report the membership inference attack (MIA) [47], which measures the MIA-Efficacy as described in [23, 13]. Finally, we assess the change in group robustness by looking at equalized odds (EO) [16], and the test accuracy of the forget-set dominant group (GA). To ease the interpretation over six metrics, we follow previous works [23, 13] and compute the average gap (Avg. Gap) and per-metric deltas with the gold standard.333Deltas are computed per each seed and then averaged over three runs. Following them, we do not report whether metrics should be maximized or minimized as the machine unlearning objective is to reduce the discrepancy with the gold standard on each metric, except for Avg. Gap which must be maximized. We use Retrain + reweight as the gold standard since it better reduces the gap with Pretrain in TA, EO, and GA, compared to Retrain and Retrain + group-DRO [42]. Further details are in the Appendix.
@@ 4.2 Results on group unlearning
| Method | reweight | RA | UA | TA | MIA | EO | GA | Avg. Gap |
| Pretrain | 84.6 | 66.6 | 82.8 | 0.1 | 20.9 | 67.3 | - | |
| Retrain | 84.7 | 54.6 | 82.3 | 0.1 | 31.6 | 56.9 | - | |
| Retrain | 81.4 | 82.8 | 79.9 | 11.3 | 3.9 | 83.5 | - | |
| Retrain | 84.5 | 66.6 | 82.4 | 0.3 | 20.8 | 67.8 | - | |
| L1-sparse [23] | 83.8 (0.7) | 44.3 (22.3) | 81.3 (1.1) | 0.2 (0.1) | 37.5 (16.7) | 47.5 (20.2) | 89.8 | |
| SalUn [13] | 84.9 (0.4) | 47.6 (19.0) | 82.4 (0.2) | 0.1 (0.2) | 31.6 (10.8) | 48.8 (19.0) | 91.7 | |
| SCRUB [28] | 82.1 (2.5) | 40.6 (26.0) | 79.8 (2.6) | 0.4 (0.2) | 40.1 (19.3) | 42.6 (25.1) | 87.4 | |
| MIU | 84.8 (0.3) | 55.3 (11.3) | 82.6 (0.3) | 0.3 (0.2) | 27.4 (6.6) | 55.9 (11.9) | 94.9 | |
| L1-sparse [23] | 83.4 (1.1) | 60.6 (6.0) | 81.3 (1.1) | 0.2 (0.2) | 27.9 (7.2) | 62.0 (5.7) | 96.5 | |
| SalUn [13] | 84.4 (0.3) | 64.9 (3.7) | 82.5 (0.2) | 0.1 (0.2) | 21.4 (1.5) | 66.2 (3.0) | 98.5 | |
| SCRUB [28] | 84.4 (0.3) | 61.6 (5.0) | 82.6 (0.3) | 0.5 (0.3) | 24.0 (3.2) | 62.9 (4.8) | 97.7 | |
| MIU | 84.2 (0.4) | 68.8 (2.6) | 82.5 (0.1) | 0.1 (0.2) | 20.2 (0.6) | 69.0 (1.3) | 99.2 |
| Method | reweight | RA | UA | TA | MIA | EO | GA | Avg. Gap |
@@
| L1-sparse [23] | 99.0 (0.2) | 59.5 (7.1) | 85.6 (1.6) | 44.0 (9.5) | 32.2 (4.1) | 48.8 (11.1) | 94.4 | |
| SalUn [13] | 100.0 (1.0) | 50.0 (9.5) | 81.8 (5.4) | 90.5 (36.9) | 38.7 (10.4) | 39.3 (12.3) | 87.4 | |
| SCRUB [28] | 98.8 (0.3) | 60.7 (10.7) | 86.9 (0.7) | 45.2 (10.7) | 31.9 (4.3) | 41.7 (9.8) | 93.9 | |
@@
| L1-sparse [23] | 98.7 (0.2) | 64.3 (11.9) | 85.0 (2.2) | 46.4 (7.1) | 30.6 (2.3) | 53.7 (8.3) | 94.7 | |
| SalUn [13] | 100.0 (1.0) | 47.6 (16.7) | 81.1 (6.1) | 91.7 (38.1) | 39.0 (10.7) | 39.0 (12.5) | 85.8 | |
| SCRUB [28] | 98.9 (0.2) | 66.7 (11.9) | 87.0 (0.6) | 44.0 (9.5) | 30.9 (3.4) | 44.3 (7.3) | 94.5 | |
@@
| L1-sparse [23] | 64.0 (2.7) | 74.1 (4.8) | 56.9 (0.5) | 0.2 (0.7) | 6.1 (0.9) | 69.4 (0.4) | 98.3 | |
| SalUn [13] | 66.3 (0.3) | 66.6 (3.0) | 55.9 (0.8) | 0.3 (0.6) | 9.0 (3.4) | 60.3 (9.3) | 97.1 | |
| SCRUB [28] | 66.9 (0.3) | 65.4 (3.9) | 56.7 (0.6) | 1.0 (0.5) | 9.9 (4.3) | 61.3 (8.3) | 97.0 | |
@@
| L1-sparse [23] | 64.4 (2.3) | 72.9 (3.6) | 56.0 (0.9) | 0.3 (0.4) | 6.1 (2.0) | 67.0 (7.1) | 97.3 | |
| SalUn [13] | 65.1 (1.5) | 69.8 (5.6) | 54.8 (1.8) | 0.3 (0.4) | 6.6 (1.7) | 63.7 (5.9) | 97.2 | |
| SCRUB [28] | 66.7 (0.2) | 73.4 (4.1) | 57.2 (0.6) | 0.7 (0.3) | 6.2 (0.7) | 70.2 (1.8) | 98.7 | |
@@
Tables 1, 2 and 3 show results for group-robust unlearning on CelebA [31], Waterbirds [42], and FairFace [24] using an unlearning ratio of .
CelebA.
@@
Retrain + reweight achieves the best trade-off between original performance preservation and unlearning, showing the best gap with TA (-0.4), EO (-0.1), and GA (+0.5). The forget accuracy (UA) equals that of Pretrain, which might seem strange at first glance. Yet, UA and GA share similar values for both Pretrain and Retrain. This suggests that the model achieves a low generalization error, as the forget accuracy aligns with that of unseen samples (from the same group) regardless of whether the forget set was part of the training data. Additionally, as CelebA [31] counts numerous images, reweight easily preserves original group accuracies, thus, recovering the original model performances and showing the same UA of Pretrain. Retraining the model with group-DRO [42] generally leads to unwanted unlearning behaviors (UA increases by 16.2) and reduced TA (-2.9). Plain Retrain, instead, suffers performance degradation in the dominant group of the forget set, as we highlight in Fig. 2. Adopting the proposed reweighting strategy leads to a better trade-off in group-robust unlearning.
@@
MIU achieves the best performance preservation (Tab. 1) by scoring the highest GA both when using (69.0%) and not using (55.9%) reweight. These results highlight that mutual information improves group performance retention. Instead, existing approaches must rely on reweight to close the gap in GA and EO with MIU as they are agnostic to group-robust machine unlearning. We also notice that using reweight increases the UA for all tested algorithms. Yet, despite the method used, coupling reweight with approximate unlearning always recovers most of the original GA. We visualize this behavior in Fig. 4, which shows the same experiment as Fig. 2 while highlighting the reweight contribution. However, as MIU is strictly designed for this task, it generally achieves higher GA and lower EO than existing approaches, scoring the best Avg. Gap.
@@
Waterbirds. Similarly to CelebA [31] experiments, Retrain + reweight achieves the best discrepancy (Tab. 2) in terms of TA (-0.5), EO (+2.1), and GA (-5.0), better preserving original group robustness. Retrain + group-DRO [42], instead, increases the UA by 4.8% above Pretrain and by 36.9% above Retrain which can negatively affect the unlearning evaluation if used as the gold standard for approximate machine unlearning. Compared to the CelebA [31] case, Retrain achieves a better unlearning-preservation trade-off. Nonetheless, Retrain + reweight is still a better candidate for group-robust unlearning, always achieving the best calibration with the original model.
@@
In Waterbirds [42], our method outperforms existing methods both with and without reweight. Yet, reweight does not provide a substantial improvement as only the forget accuracy (from a delta of 8.3 to 4.8) and the EO (from a delta of 7.5 to 4.0) get significantly enhanced. Instead, reweight substantially increases L1-sparse [23] and SCRUB [28] average UA (+4.8 and +6.0), achieving higher values than Retrain + reweight (59.5%). Nonetheless, it also improves the GA, showing better dominant forget group preservation. We argue that the subtle improvement provided by reweight is caused by the limited dataset size of Waterbirds [42]. By increasing the sampling likelihood of the few dominant forget group images left, the network overfits those few samples, limiting robustness preservation. Even without reweight, MIU outperforms all baselines, regardless of whether they use reweight, further validating our design choices.
@@
FairFace. Differently from other datasets, Retrain + group-DRO [42] struggles to preserve original model accuracy (Tab. 3) in both the dominant forget group (-13.8) and the test set (-5.8). We ascribe this behavior to numerous FairFace [24] groups that make the group-robust optimization challenging. Instead, Retrain + reweight overcomes this issue as it simply reweights group sampling probabilities to match the original training dataset statistics, achieving better GA (-1.6) and a TA (-0.5). Importantly, our experiments highlight a key advantage of reweight, which functions effectively “off the shelf” with the original model hyperparameters. Unlike Retrain + group-DRO [42], it requires no additional tuning.
@@
All methods show good alignment to Retrain + reweight, even without reweighting. TA and GA are generally preserved, with SalUn [13] scoring lowest at 55.9% (delta 0.8) and 60.3% (delta 9.3) respectively. However, methods without reweight show higher UA than Retrain, indicating partial scrubbing of the forget set. Regardless, Sec. 4.5 shows that the high UA of MIU is not caused by poor unlearning but the calibration term, which recovers the original group robustness. reweight further enhances model robustness, reflected in higher GA and lower EO. Finally, MIU + reweight and SCRUB [28] + reweight achieve the same Avg. Gap, but while our method shows a better UA and GA alignment with Retrain + reweight, SCRUB [28] obtains a better alignment in EO terms. Although reweight does not fundamentally improve the Avg. Gap, it generally enhances GA and EO, promoting performance retention.
@@
Discussion. UA strongly correlates with GA after unlearning in all three datasets. A high UA-GA correlation suggests that the forget set behaves as unseen data of the same group, indicating proper unlearning. Moreover, MIU consistently approximates Retrain+reweight better than existing methods or is on par, while reweight reliably preserves performance without drawbacks in model retraining or approximate unlearning.

### v2

@@ 4 Experiments
This section describes the experimental protocol (Sec.˜4.1) and compares MIU with multiple methods in group robust unlearning (Sec.˜4.2). We then compare existing approaches and ours when varying the unlearning ratio of the forget set dominant group (Sec.˜4.3) and when sampling the forget set from multiple groups (Sec.˜4.4). Finally, Sec.˜4.5 shows a complete ablation study of MIU’s components.
@@ 4.1 Experimental protocol
Datasets. We follow established works in group-robust optimization (Sagawa et al., 2020; Idrissi et al., 2022; Liu et al., 2021; Park et al., 2022; Park & Byun, 2024) and conduct experiments on CelebA (Liu et al., 2015) and Waterbirds (Sagawa et al., 2020) datasets, setting blonde and male as the target and protected attributes for CelebA (Liu et al., 2015), while setting waterbirds and land as the target and sensitive attributes for Waterbirds (Sagawa et al., 2020). We additionally experiment with FairFace (Karkkainen & Joo, 2021), given its numerous annotations, using age as the downstream task and randomly choosing the class 20-29 and the ethnicity afro-american as target and protected attributes, respectively. Unless stated otherwise, forget-set images are sampled from groups described by the above target-sensitive attribute pairs. Moreover, unless stated otherwise, the forget set simulates the worst-case scenario where a single group is responsible for the unlearning request, leading to a high forget distribution imbalance. The unlearning ratio is defined as the proportion of samples from that particular group that have been unlearned. After unlearning, the model must have unlearned the forget data and maintained its original robustness.
Hardware and training details. We obtain Pretrain and Retrain by fine-tuning via empirical-risk minimization a ResNet-18 (He et al., 2016) pre-trained on ImageNet (Russakovsky et al., 2015) for 30 epochs, using SGD with 0.9 momentum and weight decay. The learning rate is decayed with a cosine annealing scheduler for the entire training. We additionally warm-up the learning rate for the first two epochs using a linear scheduler. We apply standard data augmentation techniques, namely, random resized crop, random horizontal flip, and input normalization (He et al., 2016). We limited fine-tuning to 10 epochs for approximate unlearning methods, searching for the optimal configuration for the other hyperparameters. The parameter of MIU is set between 1 and 10 (see Sec.˜B.5). All experiments ran on a single A100 Nvidia GPU, using PyTorch (Paszke et al., 2019).
@@
Baselines. We compare MIU against three state-of-the-art machine unlearning approaches. L1-sparse (Jia et al., 2023) forgets sensitive information by fine-tuning the original model on the remaining set with a sparsity regularization term. SalUn (Fan et al., 2024b) proposes a saliency-based unlearning that forgets data via random labeling. SCRUB (Kurmanji et al., 2024) minimizes the divergence between the unlearned and original model on the remaining set while maximizing it on the unlearning data. Following previous works (Kurmanji et al., 2024; Jia et al., 2023), we report Pretrain, and Retrain, which are computed by fine-tuning an ImageNet (Russakovsky et al., 2015) pre-trained ResNet-18 (He et al., 2016). We also report Retrain + group-DRO (Sagawa et al., 2020) to validate the proposed Retrain + reweight.
We reimplemented all three baselines following the existing codebases. For L1-sparse (Jia et al., 2023), we perform 10 fine-tuning epochs on the remaining set with a linearly decaying regularization that follows this rule: , where is the epoch, is the total number of iterations, and the initial penalty. For SalUn (Fan et al., 2024b) we followed the “small-scale” implementation of the original codebase (i.e., the implementation for CIFAR-10 (Krizhevsky et al., 2009) and SVHN (Netzer et al., 2011)) as it is meant for datasets with few classes. We first compute the saliency mask using the gradient information from the forget set, pruning 50% of the network weights. We tune the pruned weights for 10 epochs by alternating a full pass on the forget set and a full pass on the remaining set. Also, SCRUB (Kurmanji et al., 2024) is implemented by separating forget and retaining steps, following the original code. The forget step maximizes the KL divergence between the original and the unlearned model in the forget data set. Following the original paper, we stop computing it after 3, 5, or 7 epochs. Instead, the retaining step minimizes the linear combination between the cross-entropy loss and the KL divergence between the original and unlearned model, respectively scaled by 0.99 and 0.001, as reported in (Kurmanji et al., 2024). We note that all methods use the same dataset splits; therefore, they must unlearn the same forget set.
Metrics. To evaluate unlearning and group-robustness, we rely on six different metrics. The first three are remaining (RA), forget (UA), and test (TA) accuracy. We also report the membership inference attack (MIA) (Yeom et al., 2018), which measures the MIA-Efficacy as described in (Jia et al., 2023; Fan et al., 2024b). Finally, we assess the change in group robustness by looking at equalized odds (EO) (Hardt et al., 2016), and the test accuracy of the forget-set dominant group (GA).
GA measures the ratio of correctly classified test samples that belong to the same dominant group(s) of the forget set. Therefore, let be the subset of the test set composed of all samples of group , then GA is computed as follows:
@@
Finally, to ease the interpretation of six metrics, we follow previous works (Jia et al., 2023; Fan et al., 2024b) and compute the average gap (avg. gap) and per-metric deltas with the gold standard.333Deltas are computed per each seed and then averaged over three runs. Following them, we do not report whether metrics should be maximized or minimized as the machine unlearning objective is to reduce the discrepancy with the gold standard on each metric, except for avg. gap, which must be maximized. We use Retrain + reweight as the gold standard since it better reduces the gap with Pretrain in TA, EO, and GA, compared to Retrain and Retrain + group-DRO (Sagawa et al., 2020). Further details are in Appx.˜A.
@@ 4.2 Results on group unlearning
| method | RW | RA | UA | TA | MIA | EO | GA | avg. gap |
| Pretrain | 96.2 | 41.9 | 95.9 | 0.9 | 24.2 | 40.6 | - | |
| Retrain | 96.5 | 31.3 | 95.9 | 1.6 | 27.0 | 34.4 | - | |
| Retrain + gDRO | 95.8 | 67.4 | 95.1 | 13.2 | 12.5 | 63.1 | - | |
| Retrain | 96.3 | 39.7 | 95.8 | 2.2 | 23.9 | 41.3 | - | |
| L1-sparse (Jia et al., 2023) | 95.7 (0.5) | 29.0 (10.7) | 95.4 (0.3) | 1.5 (1.3) | 28.5 (4.6) | 30.4 (10.9) | 95.3 | |
| SalUn (Fan et al., 2024b) | 96.2 (0.2) | 29.3 (10.3) | 95.8 (0.1) | 0.7 (1.5) | 29.1 (5.2) | 30.6 (10.7) | 95.3 | |
| SCRUB (Kurmanji et al., 2024) | 96.5 (0.3) | 35.1 (4.6) | 95.9 (0.2) | 0.6 (1.7) | 26.6 (2.8) | 35.9 (5.4) | 97.5 | |
| MIU | 96.4 (0.3) | 36.3 (3.4) | 95.9 (0.1) | 1.0 (1.3) | 26.1 (2.2) | 36.3 (5.0) | 98.0 | |
| L1-sparse (Jia et al., 2023) | 95.6 (0.6) | 37.3 (3.6) | 95.4 (0.4) | 0.7 (1.5) | 26.7 (2.8) | 34.8 (6.5) | 97.4 | |
| SalUn (Fan et al., 2024b) | 96.1 (0.2) | 42.9 (8.2) | 95.8 (0.1) | 0.6 (2.0) | 23.9 (3.4) | 41.1 (9.8) | 96.1 | |
| SCRUB (Kurmanji et al., 2024) | 96.4 (0.3) | 43.5 (3.8) | 96.0 (0.2) | 0.7 (1.5) | 23.7 (0.7) | 43.0 (1.7) | 98.6 | |
| MIU | 96.3 (0.3) | 43.2 (3.5) | 96.0 (0.2) | 1.2 (1.3) | 24.0 (0.5) | 41.3 (0.4) | 99.0 |
| method | RW | RA | UA | TA | MIA | EO | GA | avg. gap |
@@
| L1-sparse (Jia et al., 2023) | 99.0 (0.2) | 59.5 (7.1) | 85.6 (1.6) | 44.0 (9.5) | 32.2 (4.1) | 48.8 (11.1) | 94.4 | |
| SalUn (Fan et al., 2024b) | 100.0 (1.0) | 50.0 (9.5) | 81.8 (5.4) | 90.5 (36.9) | 38.7 (10.4) | 39.3 (12.3) | 87.4 | |
| SCRUB (Kurmanji et al., 2024) | 98.8 (0.3) | 60.7 (10.7) | 86.9 (0.7) | 45.2 (10.7) | 31.9 (4.3) | 41.7 (9.8) | 93.9 | |
@@
| L1-sparse (Jia et al., 2023) | 98.7 (0.2) | 64.3 (11.9) | 85.0 (2.2) | 46.4 (7.1) | 30.6 (2.3) | 53.7 (8.3) | 94.7 | |
| SalUn (Fan et al., 2024b) | 100.0 (1.0) | 47.6 (16.7) | 81.1 (6.1) | 91.7 (38.1) | 39.0 (10.7) | 39.0 (12.5) | 85.8 | |
| SCRUB (Kurmanji et al., 2024) | 98.9 (0.2) | 66.7 (11.9) | 87.0 (0.6) | 44.0 (9.5) | 30.9 (3.4) | 44.3 (7.3) | 94.5 | |
@@
| L1-sparse (Jia et al., 2023) | 64.0 (2.7) | 74.1 (4.8) | 56.9 (0.5) | 0.2 (0.7) | 6.1 (0.9) | 69.4 (0.4) | 98.3 | |
| SalUn (Fan et al., 2024b) | 66.3 (0.3) | 66.6 (3.0) | 55.9 (0.8) | 0.3 (0.6) | 9.0 (3.4) | 60.3 (9.3) | 97.1 | |
| SCRUB (Kurmanji et al., 2024) | 66.9 (0.3) | 65.4 (3.9) | 56.7 (0.6) | 1.0 (0.5) | 9.9 (4.3) | 61.3 (8.3) | 97.0 | |
@@
| L1-sparse (Jia et al., 2023) | 64.4 (2.3) | 72.9 (3.6) | 56.0 (0.9) | 0.3 (0.4) | 6.1 (2.0) | 67.0 (7.1) | 97.3 | |
| SalUn (Fan et al., 2024b) | 65.1 (1.5) | 69.8 (5.6) | 54.8 (1.8) | 0.3 (0.4) | 6.6 (1.7) | 63.7 (5.9) | 97.2 | |
| SCRUB (Kurmanji et al., 2024) | 66.7 (0.2) | 73.4 (4.1) | 57.2 (0.6) | 0.7 (0.3) | 6.2 (0.7) | 70.2 (1.8) | 98.7 | |
@@
Tables˜1, 2 and 3 show results for group-robust unlearning on CelebA (Liu et al., 2015), Waterbirds (Sagawa et al., 2020), and FairFace (Karkkainen & Joo, 2021) using an unlearning ratio of .
@@
CelebA. Retrain + reweight achieves the best trade-off between original performance preservation and unlearning, showing the best gap with UA (-2.2), EO (-0.3), and GA (+0.7). The forget accuracy (UA) is very close to that of Pretrain, which might seem strange at first glance. Yet, UA and GA share similar values for both Pretrain and Retrain. This suggests that the model achieves a low generalization error, as the forget accuracy aligns with that of unseen samples (from the same group) regardless of whether the forget set was part of the training data. Additionally, as CelebA (Liu et al., 2015) counts numerous images, reweight easily preserves original group accuracies, thus, recovering the original model performances and showing the same UA of Pretrain. Retraining the model with group-DRO (Sagawa et al., 2020) generally leads to unwanted unlearning behaviors, i.e., the UA increases by 25.5 points. Plain Retrain, instead, suffers performance degradation in the dominant group of the forget set, as we highlight in Fig.˜2 (GA lowers by 6.2 points). Adopting the proposed reweighting strategy leads to a better trade-off in group-robust unlearning.
@@
MIU achieves the best performance preservation (Tab.˜1) by scoring the highest GA alignment both when using (0.4) and not using (5.0) reweight. These results highlight that mutual information improves group performance retention. Instead, existing approaches must rely on reweight to close the gap in GA and EO with MIU as they are agnostic to group-robust machine unlearning. We also notice that using reweight increases the UA for all tested algorithms. Yet, despite the method used, coupling reweight with approximate unlearning always recovers most of the original GA. We visualize this in Fig.˜4, which shows the same experiment as Fig.˜2 while highlighting the reweight contribution. However, as MIU is strictly designed for this task, it generally achieves better robustness than existing approaches, scoring the best avg. gap.
@@
Waterbirds. Similarly to CelebA (Liu et al., 2015) experiments, Retrain + reweight achieves the best discrepancy (Tab.˜2) in terms of TA (-0.5), EO (+2.1), and GA (-5.0), better preserving original group robustness. Retrain + group-DRO (Sagawa et al., 2020), instead, increases the UA by 4.8% above Pretrain and by 36.9% above Retrain, which can negatively affect the unlearning evaluation if used as the gold standard for approximate machine unlearning. Compared to the CelebA (Liu et al., 2015) case, Retrain achieves a better unlearning-preservation trade-off. Nonetheless, Retrain + reweight is still a better candidate for group-robust unlearning, always achieving the best calibration with the original model.
@@
In Waterbirds (Sagawa et al., 2020), our method outperforms existing methods both with and without reweight. Yet, reweight does not provide a substantial improvement as only the forget accuracy (from a delta of 8.3 to 4.8) and the EO (from a delta of 7.5 to 4.0) get significantly enhanced. Instead, reweight substantially increases L1-sparse (Jia et al., 2023) and SCRUB (Kurmanji et al., 2024) average UA (+4.8 and +6.0), achieving higher values than Retrain + reweight (59.5%). Nonetheless, it also improves the GA, showing better dominant forget group preservation. We argue that the subtle improvement provided by reweight is caused by the limited dataset size of Waterbirds (Sagawa et al., 2020). By increasing the sampling likelihood of the few dominant forget group images left, the network overfits those few samples, limiting robustness preservation. Even without reweight, MIU outperforms all baselines, regardless of whether they use reweight, further validating our design choices.
@@
FairFace. Differently from the other datasets, here Retrain + group-DRO (Sagawa et al., 2020) struggles to preserve original model accuracy (Tab.˜3) in both the dominant forget group (-13.8) and the test set (-5.8). We ascribe this behavior to numerous FairFace (Karkkainen & Joo, 2021) groups that make the group-robust optimization challenging. Instead, Retrain + reweight overcomes this issue as it simply reweights group sampling probabilities to match the original training dataset statistics, achieving better GA (-1.6) and a TA (-0.5). Importantly, our experiments highlight a key advantage of reweight, which functions effectively “off the shelf” with the original model hyperparameters. Unlike Retrain + group-DRO (Sagawa et al., 2020), it requires no additional tuning.
@@
All methods show good alignment to Retrain + reweight, even without reweighting. TA and GA are generally preserved, with SalUn (Fan et al., 2024b) scoring lowest at 55.9% (delta 0.8) and 60.3% (delta 9.3) respectively. However, methods without reweight show higher UA than Retrain, indicating partial scrubbing of the forget set. Regardless, Sec.˜4.5 shows that the high UA of MIU is not caused by poor unlearning but the calibration term, which recovers the original group robustness. reweight further enhances model robustness, reflected in higher GA and lower EO. Finally, MIU + reweight and SCRUB (Kurmanji et al., 2024) + reweight achieve the same avg. gap, but while our method shows a better UA and GA alignment with Retrain + reweight, SCRUB (Kurmanji et al., 2024) obtains a better alignment in EO terms. Although reweight does not fundamentally improve the avg. gap, it generally enhances GA and EO, promoting performance retention.
@@
Discussion. UA strongly correlates with GA after unlearning in all three datasets. A high UA-GA correlation suggests that the forget set behaves as unseen data of the same group, indicating that the forget set was properly unlearned. Moreover, MIU consistently approximates Retrain+reweight better than existing methods or is on par, while reweight reliably preserves performance without drawbacks in model retraining or approximate unlearning.